OpenStudy (vortish):

use the system of elimination method to solve the system of equation a +8b=12 a +3b = 2

OpenStudy (anonymous):

R1=(a+8b=12) R2=(a+3b=2) R1-R2 5b=10 b=2

OpenStudy (vortish):

ok if b = 2 then of the answers give it is a ordered pair of (0,2) a solution that infinite solutions or no solutions

OpenStudy (aripotta):

you have to go back and solve for a

OpenStudy (anonymous):

back- substitution replace b=2 into R1 or R2 and solve for a

OpenStudy (vortish):

I understand that guys im just having trouble i have the subsitution for b in there just lost on what to do next

OpenStudy (aripotta):

a+8(2)=12 solve for a :)

OpenStudy (vortish):

if I understand it right a = 0

OpenStudy (aripotta):

no

OpenStudy (vortish):

no ari if your equation is right the a = -4

OpenStudy (aripotta):

yes

OpenStudy (vortish):

so the it would be a ordered pair with (2,-4) being the answer

OpenStudy (aripotta):

if it's (a,b) if would be (-4,2)

OpenStudy (aripotta):

it*

OpenStudy (vortish):

ok